### v1

@@ Abstract
The evolution of Large Language Models (LLMs) from passive text processors to autonomous agents has established planning as a core component of modern intelligence. However, achieving generalized planning remains elusive, not only by the scarcity of high-quality interaction data but also by inherent conflicts across heterogeneous planning tasks. These challenges result in models that excel at isolated tasks yet struggle to generalize, while existing multi-task training attempts suffer from gradient interference. In this paper, we present MagicAgent, a series of foundation models specifically designed for generalized agent planning. We introduce a lightweight and scalable synthetic data framework that generates high-quality trajectories across diverse planning tasks, including hierarchical task decomposition, tool-augmented planning, multi-constraint scheduling, procedural logic orchestration, and long-horizon tool execution. To mitigate training conflicts, we propose a two-stage training paradigm comprising supervised fine-tuning followed by multi-objective reinforcement learning over both static datasets and dynamic environments. Empirical results demonstrate that MagicAgent-32B and MagicAgent-30B-A3B deliver superior performance, achieving accuracies of $75.1\%$ on Worfbench, $55.9\%$ on NaturalPlan, $57.5\%$ on $\tau^{2}$-Bench, $86.9\%$ on BFCL-v3, and $81.2\%$ on ACEBench, as well as strong results on our in-house MagicEval benchmarks. These results substantially outperform existing sub-100B models and even surpass leading closed-source models.
@@ 1 INTRODUCTION
Researchers have begun to explore agent-centric model optimization to support various task-solving requirements. This paradigm typically involves fine-tuning models on task-specific trajectories to enable complex objective decomposition and task execution. A critical bottleneck, however, lies in the scarcity of high-quality training data. Prior synthetic data generation approaches have largely relied on simulated human–agent interactions within virtual environments, which are labor-intensive and cost-prohibitive (Martinez-Gonzalez et al., 2020). Furthermore, while progress has been made in tool-based planning, the broader landscape of general-purpose agent planning remains under-explored. Existing specialized models often focus on isolated capabilities, resulting in “single-point optimizations” rather than a unified planning foundation. Moreover, naive aggregation of multi-task planning data frequently leads to the “seesaw effect”, where competing optimization objectives across heterogeneous tasks impede stable joint convergence (Zhang & Yang, 2021; Yang et al., 2025a; Dong et al., 2024, 2025a). In such scenarios, performance gains in one domain often lead to unintended degradation in others, challenging the scalability of agentic models.
@@ 3.1 Hierarchical Task Decomposition
Figure 2: The synthesis pipeline of hierarchical task decomposition
@@ 3.2 Tool-Augmented Planning
Figure 3: The synthesis pipeline of tool use planning
@@ 3.3 Multi-Constraint Scheduling
Figure 4: Framework for Generating Schedule Planning Data
@@ 3.4 Procedural Logic Orchestration
Figure 5: Data construction process of Procedural Logic Orchestration
@@ 3.5 Long-Horizon Tool Execution
Figure 6: The pipeline to synthesis long-horizon tool execution trajectory data.
@@ 6.3 Performance of Online RL
(b) Online and offline RL rewards
@@ 6.4 Performance of MoE
(a) Calendar task
(b) Meeting task
(c) Trip task
@@ Appendix A Training Details
We provide a comprehensive overview of the training hyperparameters for MagicAgent-32B and MagicAgent-30B-A3B, including core model configurations, mixture-of-experts (MoE) parallelization strategies, and optimization settings. Table 8 details prompt and generation limits, optimizer and scheduler choices, and numerical precision, along with parallelism parameters tailored for large-scale MoE training. It further specifies the hyperparameters used for both Supervised Fine-tuning (SFT) and Reinforcement Learning (RL), including batch sizes, learning rates, generation strategies, and reward weighting, ensuring reproducibility and stable optimization across training stages.
@@
max_prompt_length 20480 Maximum prompt length
max_new_token 2048 Maximum completion length
@@
num_gpus 32 Number of GPUs (4 machines × 8 GPUs)
@@
num_gpus 32 Number of GPUs (4 machines × 8 GPUs)
@@ Appendix C Hierarchical Task Decomposition Examples
Raw Format <Query> åÿőæĹŚåIJĺç¿ŐåŻćéćĎèőćæİŋåůđèě£æźŰçŽĎéĚŠåžŮïijŇåĚěä¡Ŕ11æIJĹ5æŮěïijŇçężåžŮ11æIJĹ7æŮěïijŇèęĄèőćäÿďäÿłæĹ£éŮťïijŇéąžä¿£æ§ěäÿĂæ§ěéŹĎè£Śåě¡åŘČçŽĎéďŘéęĘãĂĆ </Query> <Plan> Step1. åIJĺç¿ŐåŻćäÿŋéćĎèőćä¡ŊäžŐæİŋåůđçŽĎèě£æźŰéĚŠåžŮïijŇåĚěä¡ŔæŮěæIJ§äÿž11æIJĹ5æŮěïijŇçężåžŮæŮěæIJ§äÿž11æIJĹ7æŮěïijŇéćĎåőŽ2äÿłæĹ£éŮť Step2. åIJĺç¿ŐåŻćäÿŋæŘIJçťćèě£æźŰéĚŠåžŮéŹĎè£ŚçŽĎç¿Őéč§éďŘéęĘïijŇåźűæŇĽåě¡èŕĎæŐŠåžŔ </Plan>

### v2

@@ Abstract
The evolution of Large Language Models (LLMs) from passive text processors to autonomous agents has established planning as a core component of modern intelligence. However, achieving generalized planning remains elusive, not only by the scarcity of high-quality interaction data but also by inherent conflicts across heterogeneous planning tasks. These challenges result in models that excel at isolated tasks yet struggle to generalize, while existing multi-task training attempts suffer from gradient interference. In this paper, we present MagicAgent, a series of foundation models specifically designed for generalized agent planning. We introduce a lightweight and scalable synthetic data framework that generates high-quality trajectories across diverse planning tasks, including hierarchical task decomposition, tool-augmented planning, multi-constraint scheduling, procedural logic orchestration, and long-horizon tool execution. To mitigate training conflicts, we propose a two-stage training paradigm comprising supervised fine-tuning followed by multi-objective reinforcement learning over both static datasets and dynamic environments. Empirical results show that MagicAgent-32B and MagicAgent-30B-A3B achieve superior performance across diverse open-source benchmarks (e.g., $75.1\%$ on Worfbench and $86.9\%$ on BFCL-v3), as well as strong results on our in-house MagicEval benchmarks, substantially outperforming existing sub-100B models and surpassing leading ultra-scale models, including GPT-5.2, Kimi-K2 and GLM-4.7.
@@ 1 INTRODUCTION
Researchers have increasingly explored agent-centric model optimization to support diverse task-solving requirements. This paradigm typically involves fine-tuning models on task-specific trajectories to facilitate complex objective decomposition and execution. However, the development of robust, general-purpose agent planning is currently hindered by three primary limitations in existing methodologies: (1) Data Scarcity and Generation Overhead: A fundamental bottleneck lies in the scarcity of high-quality training data. Prior synthetic data generation approaches have largely relied on simulated human-agent interactions within virtual environments—a process that is both labor-intensive and cost-prohibitive (Martinez-Gonzalez et al., 2020; Huang et al., 2024). (2) Fragmented Capabilities in Bespoke Systems: While progress has been made in tool-based planning, the broader landscape of general-purpose agent planning remains under-explored (Wang et al., 2025; Yuan et al., 2025). Existing bespoke systems frequently focus on isolated capabilities; such single-point optimizations inherently lack the generalization ability and scalability required to establish a unified planning foundation. (3) Optimization Instability in Multi-Task Learning: Naive aggregation of multi-task planning data frequently induces a “seesaw effect”. Competing optimization objectives across heterogeneous tasks impede stable joint convergence (Zhang & Yang, 2021; Yang et al., 2025a; Dong et al., 2024, 2025a). In such scenarios, performance gains in one domain often trigger unintended degradation in others, further challenging the scalability of current agentic models.
@@ 3.1 Hierarchical Task Decomposition
Figure 2: The synthesis pipeline of hierarchical task decomposition.
@@ 3.2 Tool-Augmented Planning
Figure 3: The synthesis pipeline of tool use planning.
@@ 3.3 Multi-Constraint Scheduling
Figure 4: The synthesis pipeline of multi-constrain scheduling.
@@ 3.4 Procedural Logic Orchestration
Figure 5: The synthesis pipeline of procedural logic orchestration.
@@ 3.5 Long-Horizon Tool Execution
Figure 6: The synthesis pipeline of long-horizon tool execution.
@@ 6.3 Performance of Online RL
(b) Online and offline RL rewards.
@@ 6.4 Performance of MoE
(a) Calendar task.
(b) Meeting task.
(c) Trip task.
@@ Appendix A Training Details
We provide a comprehensive overview of the training hyperparameters for MagicAgent-32B and MagicAgent-30B-A3B-Instruct-2507, including core model configurations, mixture-of-experts (MoE) parallelization strategies, and optimization settings. Table 8 details prompt and generation limits, optimizer and scheduler choices, and numerical precision, along with parallelism parameters tailored for large-scale MoE training. It further specifies the hyperparameters used for both Supervised Fine-tuning (SFT) and Reinforcement Learning (RL), including batch sizes, learning rates, generation strategies, and reward weighting, ensuring reproducibility and stable optimization across training stages.
@@
max_length 32768 Maximum prompt length and completion length
@@
num_gpus 32 Number of GPUs (4 machines $\times$ 8 GPUs)
@@
num_gpus 32 Number of GPUs (4 machines $\times$ 8 GPUs)
@@ Appendix C Hierarchical Task Decomposition Examples
Raw Format <Query> Help me book a hotel at West Lake in Hangzhou on Meituan, checking in on November 5th and checking out on November 7th. I need to book two rooms. Also, check for good restaurants nearby. </Query> <Plan> Step1. Book a hotel at West Lake in Hangzhou on Meituan, with a check-in date of November 5th and a check-out date of November 7th, booking 2 rooms Step2. Search for gourmet restaurants near the West Lake hotel on Meituan and sort by high ratings </Plan>